# Fermion and Boson Pairs in Beamsplitters and MZIs

### v1

@@ III.2 Adding Distinguishability
$\begin{split}\left|\psi\right\rangle&=\frac{1}{2}\left(\sqrt{r_{1,2}}\left((\hat{c}^{\dagger}_{1})^{2}-\hat{c}^{\dagger}_{1}\hat{d}^{\dagger}_{1}+\hat{d}^{\dagger}_{1}\hat{c}^{\dagger}_{1}-(\hat{d}^{\dagger}_{1})^{2}\right)+\sqrt{1-r_{1,2}}\left(\hat{c}^{\dagger}_{1}\hat{c}^{\dagger}_{1\perp}-\hat{c}^{\dagger}_{1}\hat{d}^{\dagger}_{1\perp}+\hat{d}^{\dagger}_{1}\hat{c}^{\dagger}_{1\perp}-\hat{d}^{\dagger}_{1}\hat{d}^{\dagger}_{1\perp}\right)\right)\left|0,0\right\rangle\end{split}$ (25)
@@
$\begin{split}\left|\psi\right\rangle&=\frac{1}{2}\left(\sqrt{r_{1,2}}\left((\hat{a}^{\dagger})^{2}-\hat{a}^{\dagger}\hat{b}^{\dagger}+\hat{b}^{\dagger}\hat{a}^{\dagger}-(\hat{b}^{\dagger})^{2}\right)+\sqrt{1-r_{1,2}}\left(\hat{a}^{\dagger}\hat{a}^{\dagger}_{\perp}-\hat{a}^{\dagger}\hat{b}^{\dagger}_{\perp}+\hat{b}^{\dagger}\hat{a}^{\dagger}_{\perp}-\hat{b}^{\dagger}\hat{b}^{\dagger}_{\perp}\right)\right)\left|0,0\right\rangle\\ &=\left(\sqrt{r_{1,2}}\hat{c}^{\dagger}\hat{d}^{\dagger}+\sqrt{1-r_{1,2}}\hat{c}^{\dagger}\hat{d}^{\dagger}_{\perp}\right)\left|0,0\right\rangle=\left|1,1\right\rangle\end{split}$ (26)
@@
$\begin{split}\left|\psi\right\rangle&=\frac{1}{2}\Bigg{(}\sqrt{r_{1,2}}\left((\hat{a}^{\dagger})^{2}-e^{i\phi}\hat{a}^{\dagger}\hat{b}^{\dagger}+e^{i\phi}\hat{b}^{\dagger}\hat{a}^{\dagger}-e^{2i\phi}(\hat{b}^{\dagger})^{2}\right)\\ &\;\;+\sqrt{1-r_{1,2}}\left(\hat{a}^{\dagger}\hat{a}^{\dagger}_{\perp}-e^{i\phi}\hat{a}^{\dagger}\hat{b}^{\dagger}_{\perp}+e^{i\phi}\hat{b}^{\dagger}\hat{a}^{\dagger}_{\perp}-e^{2i\phi}\hat{b}^{\dagger}\hat{b}^{\dagger}_{\perp}\right)\Bigg{)}\left|0,0\right\rangle\\ &=\frac{1}{2}\Bigg{(}\frac{\sqrt{r_{1,2}}}{2}\left((1-e^{2i\phi})(\hat{c}^{\dagger})^{2}+(1+e^{i\phi})^{2}\hat{c}^{\dagger}\hat{d}^{\dagger}+(1-e^{i\phi})^{2}\hat{d}^{\dagger}\hat{c}^{\dagger}+(1-e^{2i\phi})(\hat{d}^{\dagger})^{2}\right)\\ &\;\;+\frac{\sqrt{1-r_{1,2}}}{2}\left((1-e^{2i\phi})\hat{c}^{\dagger}\hat{c}^{\dagger}_{\perp}+(1-e^{i\phi})^{2}\hat{d}^{\dagger}\hat{c}^{\dagger}_{\perp}+(1+e^{2i\phi})^{2}\hat{c}^{\dagger}\hat{d}^{\dagger}_{\perp}+(1-e^{2i\phi})\hat{d}^{\dagger}\hat{d}^{\dagger}_{\perp}\right)\Bigg{)}\left|0,0\right\rangle\\ \end{split}$ (27)
@@ III.3 Two Bosons entering an MZI via the same Port
$\begin{split}\left|\psi\right\rangle&=\frac{1}{2\sqrt{2}}\left((\hat{a}^{\dagger})^{2}+\hat{a}^{\dagger}\hat{b}^{\dagger}+\hat{b}^{\dagger}\hat{a}^{\dagger}+(\hat{b}^{\dagger})^{2}\right)\left|0,0\right\rangle\\ &=\frac{1}{4\sqrt{2}}\left((\hat{c}^{\dagger}+\hat{d}^{\dagger})^{2}+(\hat{c}^{\dagger}+\hat{d}^{\dagger})(\hat{c}^{\dagger}-\hat{d}^{\dagger})+(\hat{c}^{\dagger}-\hat{d}^{\dagger})(\hat{c}^{\dagger}+\hat{d}^{\dagger})+(\hat{c}^{\dagger}-\hat{d}^{\dagger})^{2}\right)\left|0,0\right\rangle\\ &=\frac{1}{4\sqrt{2}}\left(4\hat{c}^{\dagger}\hat{c}^{\dagger}\right)\left|0,0\right\rangle=\frac{1}{\sqrt{2}}\hat{c}^{\dagger}\hat{c}^{\dagger}\left|0,0\right\rangle=\left|2,0\right\rangle\end{split}$ (29)
@@
$\begin{split}\left|\psi\right\rangle&=\frac{1}{2\sqrt{2}}\left((\hat{a}^{\dagger})^{2}+e^{i\phi}\hat{a}^{\dagger}\hat{b}^{\dagger}+e^{i\phi}\hat{b}^{\dagger}\hat{a}^{\dagger}+e^{2i\phi}(\hat{b}^{\dagger})^{2}\right)\left|0,0\right\rangle\\ &=\frac{1}{4\sqrt{2}}\left((\hat{c}^{\dagger}+\hat{d}^{\dagger})^{2}+e^{i\phi}(\hat{c}^{\dagger}+\hat{d}^{\dagger})(\hat{c}^{\dagger}-\hat{d}^{\dagger})+e^{i\phi}(\hat{c}^{\dagger}-\hat{d}^{\dagger})(\hat{c}^{\dagger}+\hat{d}^{\dagger})+e^{2i\phi}(\hat{c}^{\dagger}-\hat{d}^{\dagger})^{2}\right)\left|0,0\right\rangle\\ &=\frac{1}{4\sqrt{2}}\left((1+e^{i\phi})^{2}\hat{c}^{\dagger}\hat{c}^{\dagger}+(1-e^{2i\phi})\hat{d}^{\dagger}\hat{c}^{\dagger}+(1-e^{2i\phi})\hat{c}^{\dagger}\hat{d}^{\dagger}+(1-e^{i\phi})^{2}\hat{d}^{\dagger}\hat{d}^{\dagger}\right)\left|0,0\right\rangle\\ &=\frac{(1+e^{i\phi})^{2}}{4}\left|2,0\right\rangle+\frac{1-e^{2i\phi}}{2\sqrt{2}}\left|1,1\right\rangle+\frac{(1-e^{i\phi})^{2}}{4}\left|0,2\right\rangle\end{split}$ (30)
@@
$\begin{split}\left|\psi\right\rangle&=\frac{1}{2\sqrt{2}}\Bigg{(}\sqrt{r_{1,2}}\left((\hat{a}^{\dagger})^{2}+e^{i\phi}\hat{a}^{\dagger}\hat{b}^{\dagger}+e^{i\phi}\hat{b}^{\dagger}\hat{a}^{\dagger}+e^{2i\phi}(\hat{b}^{\dagger})^{2}\right)\\ &\;\;+\sqrt{1-r_{1,2}}\left(\hat{a}^{\dagger}\hat{a}^{\dagger}_{\perp}+e^{i\phi}\hat{a}^{\dagger}\hat{b}^{\dagger}_{\perp}+e^{i\phi}\hat{b}^{\dagger}\hat{a}^{\dagger}_{\perp}+e^{2i\phi}\hat{b}^{\dagger}\hat{b}^{\dagger}_{\perp}\right)\Bigg{)}\left|0,0\right\rangle\\ &=\Bigg{(}\frac{\sqrt{r_{1,2}}}{4\sqrt{2}}\left((1+e^{i\phi})^{2}\hat{c}^{\dagger}\hat{c}^{\dagger}+(1-e^{2i\phi})(\hat{d}^{\dagger}\hat{c}^{\dagger}+\hat{c}^{\dagger}\hat{d}^{\dagger})+(1-e^{i\phi})^{2}\hat{d}^{\dagger}\hat{d}^{\dagger}\right)\\ &\;\;+\frac{\sqrt{1-r_{1,2}}}{4\sqrt{2}}\left((1+e^{i\phi})^{2}\hat{c}^{\dagger}\hat{c}^{\dagger}_{\perp}+(1-e^{2i\phi})(\hat{d}^{\dagger}\hat{c}^{\dagger}_{\perp}+\hat{c}^{\dagger}\hat{d}^{\dagger}_{\perp})+(1-e^{i\phi})^{2}\hat{d}^{\dagger}\hat{d}^{\dagger}_{\perp}\right)\Bigg{)}\left|0,0\right\rangle\\ &=\frac{1}{4\sqrt{2}}\Bigg{(}(1+e^{i\phi})^{2}\hat{c}^{\dagger}_{1}\hat{c}^{\dagger}_{2}+(1-e^{2i\phi})(\hat{d}^{\dagger}_{1}\hat{c}^{\dagger}_{2}+\hat{c}^{\dagger}_{1}\hat{d}^{\dagger}_{2})+(1-e^{i\phi})^{2}\hat{d}^{\dagger}_{1}\hat{d}^{\dagger}_{2}\Bigg{)}\left|0,0\right\rangle\end{split}$ (31)
mimicking the result above, regardless of how distinguishable the bosons are—at $\phi=0$, both bosons come out of port $c$; at $\phi=\pi$, both bosons come out of port $d$; and at $\phi=\pi/2$, there’s a 50% chance the two photons come out of opposite ports, and a 25% chance each that both photons come out of port $c$ and port $d$ respectively.
@@
IV Summary
@@
Acknowledgements - We thank Mike Taverne for useful comments.

### v2

@@ III.2 Adding Distinguishability
$\begin{split}\left|\psi\right\rangle&=\frac{-1}{2}\left(\sqrt{r_{1,2}}\left((\hat{c}^{\dagger}_{1})^{2}-\hat{c}^{\dagger}_{1}\hat{d}^{\dagger}_{1}+\hat{d}^{\dagger}_{1}\hat{c}^{\dagger}_{1}-(\hat{d}^{\dagger}_{1})^{2}\right)+\sqrt{1-r_{1,2}}\left(\hat{c}^{\dagger}_{1}\hat{c}^{\dagger}_{1\perp}-\hat{c}^{\dagger}_{1}\hat{d}^{\dagger}_{1\perp}+\hat{d}^{\dagger}_{1}\hat{c}^{\dagger}_{1\perp}-\hat{d}^{\dagger}_{1}\hat{d}^{\dagger}_{1\perp}\right)\right)\left|0,0\right\rangle\end{split}$ (26)
@@
$\begin{split}\left|\psi\right\rangle&=\frac{-1}{2}\left(\sqrt{r_{1,2}}\left((\hat{a}^{\dagger})^{2}-\hat{a}^{\dagger}\hat{b}^{\dagger}+\hat{b}^{\dagger}\hat{a}^{\dagger}-(\hat{b}^{\dagger})^{2}\right)+\sqrt{1-r_{1,2}}\left(\hat{a}^{\dagger}\hat{a}^{\dagger}_{\perp}-\hat{a}^{\dagger}\hat{b}^{\dagger}_{\perp}+\hat{b}^{\dagger}\hat{a}^{\dagger}_{\perp}-\hat{b}^{\dagger}\hat{b}^{\dagger}_{\perp}\right)\right)\left|0,0\right\rangle\\ &=-\left(\sqrt{r_{1,2}}\hat{d}^{\dagger}\hat{c}^{\dagger}+\sqrt{1-r_{1,2}}\hat{d}^{\dagger}\hat{c}^{\dagger}_{\perp}\right)\left|0,0\right\rangle=-\left|1,1\right\rangle\end{split}$ (27)
@@
$\begin{split}\left|\psi\right\rangle&=\frac{-1}{2}\Bigg{(}\sqrt{r_{1,2}}\left((\hat{a}^{\dagger})^{2}-e^{i\phi}\hat{a}^{\dagger}\hat{b}^{\dagger}+e^{i\phi}\hat{b}^{\dagger}\hat{a}^{\dagger}-e^{2i\phi}(\hat{b}^{\dagger})^{2}\right)\\ &\;\;+\sqrt{1-r_{1,2}}\left(\hat{a}^{\dagger}\hat{a}^{\dagger}_{\perp}-e^{i\phi}\hat{a}^{\dagger}\hat{b}^{\dagger}_{\perp}+e^{i\phi}\hat{b}^{\dagger}\hat{a}^{\dagger}_{\perp}-e^{2i\phi}\hat{b}^{\dagger}\hat{b}^{\dagger}_{\perp}\right)\Bigg{)}\left|0,0\right\rangle\\ &=\frac{-1}{2}\Bigg{(}\frac{\sqrt{r_{1,2}}}{2}\left((1-e^{2i\phi})(\hat{c}^{\dagger})^{2}+(1-e^{i\phi})^{2}\hat{c}^{\dagger}\hat{d}^{\dagger}+(1+e^{i\phi})^{2}\hat{d}^{\dagger}\hat{c}^{\dagger}+(1-e^{2i\phi})(\hat{d}^{\dagger})^{2}\right)\\ &\;\;+\frac{\sqrt{1-r_{1,2}}}{2}\left((1-e^{2i\phi})\hat{c}^{\dagger}\hat{c}^{\dagger}_{\perp}+(1-e^{2i\phi})^{2}\hat{c}^{\dagger}\hat{d}^{\dagger}_{\perp}+(1+e^{i\phi})^{2}\hat{d}^{\dagger}\hat{c}^{\dagger}_{\perp}+(1-e^{2i\phi})\hat{d}^{\dagger}\hat{d}^{\dagger}_{\perp}\right)\Bigg{)}\left|0,0\right\rangle\\ \end{split}$ (28)
@@ III.3 Two Bosons entering an MZI via the same Port
$\begin{split}\left|\psi\right\rangle&=\frac{1}{2\sqrt{2}}\left((\hat{a}^{\dagger})^{2}+\hat{a}^{\dagger}\hat{b}^{\dagger}+\hat{b}^{\dagger}\hat{a}^{\dagger}+(\hat{b}^{\dagger})^{2}\right)\left|0,0\right\rangle\\ &=\frac{1}{4\sqrt{2}}\left((\hat{c}^{\dagger}+\hat{d}^{\dagger})^{2}-(\hat{c}^{\dagger}+\hat{d}^{\dagger})(\hat{c}^{\dagger}-\hat{d}^{\dagger})-(\hat{c}^{\dagger}-\hat{d}^{\dagger})(\hat{c}^{\dagger}+\hat{d}^{\dagger})+(\hat{c}^{\dagger}-\hat{d}^{\dagger})^{2}\right)\left|0,0\right\rangle\\ &=\frac{1}{4\sqrt{2}}\left(4\hat{c}^{\dagger}\hat{c}^{\dagger}\right)\left|0,0\right\rangle=\frac{1}{\sqrt{2}}\hat{c}^{\dagger}\hat{c}^{\dagger}\left|0,0\right\rangle=\left|2,0\right\rangle\end{split}$ (30)
@@
$\begin{split}\left|\psi\right\rangle&=\frac{1}{2\sqrt{2}}\left((\hat{a}^{\dagger})^{2}+e^{i\phi}\hat{a}^{\dagger}\hat{b}^{\dagger}+e^{i\phi}\hat{b}^{\dagger}\hat{a}^{\dagger}+e^{2i\phi}(\hat{b}^{\dagger})^{2}\right)\left|0,0\right\rangle\\ &=\frac{1}{4\sqrt{2}}\left((\hat{c}^{\dagger}+\hat{d}^{\dagger})^{2}-e^{i\phi}(\hat{c}^{\dagger}+\hat{d}^{\dagger})(\hat{c}^{\dagger}-\hat{d}^{\dagger})-e^{i\phi}(\hat{c}^{\dagger}-\hat{d}^{\dagger})(\hat{c}^{\dagger}+\hat{d}^{\dagger})+e^{2i\phi}(\hat{c}^{\dagger}-\hat{d}^{\dagger})^{2}\right)\left|0,0\right\rangle\\ &=\frac{1}{4\sqrt{2}}\left((1-e^{i\phi})^{2}\hat{c}^{\dagger}\hat{c}^{\dagger}+(1-e^{2i\phi})\hat{d}^{\dagger}\hat{c}^{\dagger}+(1-e^{2i\phi})\hat{c}^{\dagger}\hat{d}^{\dagger}+(1+e^{i\phi})^{2}\hat{d}^{\dagger}\hat{d}^{\dagger}\right)\left|0,0\right\rangle\\ &=\frac{(1-e^{i\phi})^{2}}{4}\left|2,0\right\rangle+\frac{1-e^{2i\phi}}{2\sqrt{2}}\left|1,1\right\rangle+\frac{(1+e^{i\phi})^{2}}{4}\left|0,2\right\rangle\end{split}$ (31)
@@
$\begin{split}\left|\psi\right\rangle&=\frac{1}{2\sqrt{2}}\Bigg{(}\sqrt{r_{1,2}}\left((\hat{a}^{\dagger})^{2}+e^{i\phi}\hat{a}^{\dagger}\hat{b}^{\dagger}+e^{i\phi}\hat{b}^{\dagger}\hat{a}^{\dagger}+e^{2i\phi}(\hat{b}^{\dagger})^{2}\right)\\ &\;\;+\sqrt{1-r_{1,2}}\left(\hat{a}^{\dagger}\hat{a}^{\dagger}_{\perp}+e^{i\phi}\hat{a}^{\dagger}\hat{b}^{\dagger}_{\perp}+e^{i\phi}\hat{b}^{\dagger}\hat{a}^{\dagger}_{\perp}+e^{2i\phi}\hat{b}^{\dagger}\hat{b}^{\dagger}_{\perp}\right)\Bigg{)}\left|0,0\right\rangle\\ &=\Bigg{(}\frac{\sqrt{r_{1,2}}}{4\sqrt{2}}\left((1-e^{i\phi})^{2}\hat{c}^{\dagger}\hat{c}^{\dagger}+(1-e^{2i\phi})(\hat{d}^{\dagger}\hat{c}^{\dagger}+\hat{c}^{\dagger}\hat{d}^{\dagger})+(1+e^{i\phi})^{2}\hat{d}^{\dagger}\hat{d}^{\dagger}\right)\\ &\;\;+\frac{\sqrt{1-r_{1,2}}}{4\sqrt{2}}\left((1-e^{i\phi})^{2}\hat{c}^{\dagger}\hat{c}^{\dagger}_{\perp}+(1-e^{2i\phi})(\hat{d}^{\dagger}\hat{c}^{\dagger}_{\perp}+\hat{c}^{\dagger}\hat{d}^{\dagger}_{\perp})+(1+e^{i\phi})^{2}\hat{d}^{\dagger}\hat{d}^{\dagger}_{\perp}\right)\Bigg{)}\left|0,0\right\rangle\\ &=\frac{1}{4\sqrt{2}}\Bigg{(}(1-e^{i\phi})^{2}\hat{c}^{\dagger}_{1}\hat{c}^{\dagger}_{2}+(1-e^{2i\phi})(\hat{d}^{\dagger}_{1}\hat{c}^{\dagger}_{2}+\hat{c}^{\dagger}_{1}\hat{d}^{\dagger}_{2})+(1+e^{i\phi})^{2}\hat{d}^{\dagger}_{1}\hat{d}^{\dagger}_{2}\Bigg{)}\left|0,0\right\rangle\end{split}$ (32)
mimicking the result above, regardless of how distinguishable the bosons are—at $\phi=0$, both bosons come out of port $d$; at $\phi=\pi$, both bosons come out of port $c$; and at $\phi=\pi/2$, there’s a 50% chance the two photons come out of opposite ports, and a 25% chance each that both photons come out of port $c$ and port $d$ respectively.
@@
IV Discussion
One thing to note is that we would expect there to be a classical correspondence principle [9] in place in this scenario: as we scale up the number of particles, especially in the case of bosons like photons, we would expect their collective behaviour to mirror that of the classical field they correspond to (e.g., light for photons). Considering the classical analogue of our indistinguishable photon pair—a pair of identical coherent plane waves, we obviously don’t see bunching when these arrive at the two ports of a beam splitter. However, analogous interference effects still occur in this scenario to those which remove the terms for the two indistinguishable photons where one photon exits from one port, and the other photon from the other port—we just see the effective many-pair average of both photons exiting from one port and both exiting from the other. These same analogous interference effects are why, in the classical case, we can balance a Mach-Zehnder interferometer such that the (coherent) light we input from one port, always exits from one port, in analogy to the result shown in Section III.3. It is interesting to note that our intuition therefore more easily accepts effects visible when we take an average of many quantum cases, but struggles with the fluctuations from this average which we only see when we look at few-particle cases. This makes sense, given we only ever see phenomena which exist at the many-particle limit in our day-to-day life, and our intuition, to some degree, must be based on what we repeatedly observe. Obviously, given we cannot scale up the number of fermions in the same mode, there is no equivalent correspondence principle for these—this may be why our intuition struggles more with the effects of fermions than bosons, unless we try and use inherently artificial classical analogies to represent these quantum effects (e.g., thinking about Pauli exclusion as if through electrostatic repulsion, or the particle “taking up space”).
Similarly, given the simplicity of the devices we consider, one may wonder why it is that the effects we give are so rarely observable, and so why we struggle to get an intuition for them. The (hopefully obvious) answer to this is that the described experiments all require us to be dealing with single (or few) particles, and for these particles to be indistinguishable (or nearly indistinguishable). Accessing single particles (e.g., through generating paired photons through spontaneous parametric down conversion or spontaneous four-wave mixing) is one of the most difficult parts of modern quantum optics/photonics experiments, and the time it took for techniques to be developed to allow us to do this are a big part of the reason it took so long for Bell’s Theorem to be proven [10, 11, 12, 13]. Despite the underpinning theory describing the behaviour of these single particles having been put down nearly a hundred years ago, we are only now arriving at the stage where these particles can be manipulated individually for technological development (the “Second Quantum Revolution” [14]). Given we have only recently gained such access to individual quantum particles, it is deeply interesting to think what we could learn about quantum mechanics by studying how they interfere, and how we could deploy such quantum interference to develop new, inherently quantum, technologies.
An interesting paper [15] has recently proposed effectively doing the reverse of what we discuss in this paper—identifying, for (the many-repeat limit of) two particles each input into one port of a Mach-Zehnder interferometer, whether the particles were bosons or fermions, and quantify how distinguishable the particles were from one another. The protocol does this by taking the output probabilities for each of the two output ports as a function of some control parameter (in the MZI case, the phase applied between the two paths), and using this to calculate the (classical) Fisher Information [16]. They show that the maximum of this Fisher Information with respect to the phase is proportional to a parameter, $\Gamma$ which is equal to 1 when the two particles are indistinguishable bosons, is equal to -1 when the two particles are indistinguishable fermions, and is goes between 1 and 0 (-1 and 0) as the bosons (fermions) become more distinguishable. While taking the opposite route to us (we pedagogically describe (single-shot) output statistics for given particle statistics, while they give a protocol to identify particle statistics from (a large sample set of) output statistics, the paper poses a useful example of the kinds of metrological protocols which can come from the quantum interference we discuss in this paper.
@@ IV Discussion
Acknowledgements - I thank Mike Taverne for useful comments.